### v1

# Contextualizing Recommendation Explanations with LLMs: A User Study

Abstract
Large language models (LLMs) are increasingly prevalent in recommender systems, where LLMs can be used to generate personalized recommendations. Here, we examine how different LLM-generated explanations for movie recommendations affect users’ perceptions of cognitive, affective, and utilitarian needs and consumption intentions. In a pre-registered, between-subject online experiment (N=759) and follow-up interviews (N=30), we compare (a) LLM-generated generic explanations, and (b) LLM-generated contextualized explanations. Our findings show that contextualized explanations (i.e., explanations that incorporate users’ past behaviors) effectively meet users’ cognitive needs while increasing users’ intentions to watch recommended movies. However, adding explanations offers limited benefits in meeting users’ utilitarian and affective needs, raising concerns about the proper design and implications of LLM-generated explanations. Qualitative insights from interviews reveal that referencing users’ past preferences enhances trust and understanding but can feel excessive if overused. Furthermore, users with more active and positive engagement with the recommender system and movie-watching get substantial gains from contextualized explanations. Overall, our research clarifies how LLM-generated recommendations influence users’ motivations and behaviors, providing valuable insights for the future development of user-centric recommender systems, a key element in social media platforms and online ecosystems.
Codehttps://anonymous.4open.science/r/Contextualizing-Recommendation-Explanations-With-LLMs-55D3
1 Introduction
Recommender systems influence human decision-making in daily life, such as purchasing products on e-commerce platforms and choosing movies on streaming services (Angelov et al. 2021; Goyani and Chaurasiya 2020). However, users often fail to understand why they receive specific recommendations. On streaming platforms like Netflix and Disney+, users are presented with recommended movies without any explanation of why those movies might be of interest. Without such explanations, users may find it difficult to judge whether the recommendation is helpful and decide on their consumption. Explanations help users make choices that reflect their preferences and improve decision quality (Vultureanu-Albisi and Badica 2021). As a result, there is a growing research interest in offering explanations behind recommendations (Ko et al. 2022; Zhang and Chen 2020).
The emergence of large language models (LLMs) presents potential for improving these recommendation explanations (Wu et al. 2024; Hou, Leach, and Huang 2024). Some studies have adopted LLMs to produce more natural and human-like explanations. However, current efforts often rely on generic explanation templates (Li, Zhang, and Chen 2020). These generic explanations overlook users’ contexts or preferences, thereby limiting the benefits of the explanations and underutilizing the LLMs’ reasoning capabilities. Existing research seldom provides a comprehensive understanding of how users perceive the received recommendations. Current work also rarely examines whether positive user perceptions translate into users’ consumption intentions.
Motivated by this gap, our study adopts established explanation measurement standards (Balog and Radlinski 2020) and interprets user responses through the lens of the Uses and Gratifications (U&G) framework (Katz, Blumler, and Gurevitch 1973; Lin et al. 2013). The U&G framework categorizes users’ diverse needs, including the desire for understanding and knowledge acquisition (Cognitive needs), the need for emotional connection and engagement (Affective needs), and the focus on practicality and task-oriented goals (Utilitarian needs). This framework provides a valuable perspective to understand how users actively select, consume, and engage with the received information. Our goal is to examine how different LLM-generated recommendation explanation types resonate with specific user needs and, in turn, influence their consumption intentions. We design the following research question:
How do different LLM-generated explanations respond to users’ diverse motivational needs (cognitive, affective, utilitarian) and affect their consumption intentions?
Figure 1: Overview of the study process.
In this study, we conducted a pre-registered, between-subject online experiment involving $N=759$ subjects from the U.S. in the context of movie recommendations. [^1] The experimental protocol was pre-registered at OSF: https://osf.io/j46vx/?view_only=dc2b1f641b33407ab8168736a783055a By drawing upon the U&G framework, we analyze users’ perceptions and intentions towards three types of recommendations generated by GPT-4 Turbo: (1) recommendations with no explanations (NE), (2) recommendations with LLM-generated generic explanations (GE), and (3) recommendations with LLM-generated contextualized explanations (CE). We also drew upon user profile information and conducted follow-up interviews ($N=30$) to interpret our results more deeply. To ensure the robustness of our study design and the generalizability of the results, we conducted supplementary experiments by using a different LLM (Claude 3.5 Sonnet) with an additional 140 users from the U.S. Figure 1 provides an overview of our study flow.
Contributions: This paper demonstrates how LLM-generated, contextualized explanations can more effectively address users’ diverse motivations and increase their consumption intentions. By integrating the U&G framework, our study shows how different explanations affect users’ cognitive, affective, and utilitarian needs. Moreover, we highlight the amplifying role of user engagement with the recommender system, revealing that users who already find the system beneficial gain the most from explanations. These findings bridge advancements in LLMs with user behavior and provide actionable insights to enhance user-centric recommender systems in social media contexts.
2 Related Work
2.1 Explanations in Recommender Systems
In social media and online communities, recommendations help users to find items aligned with their preferences. Initial research centered on improving accuracy (Ma et al. 2023; Wu et al. 2023), but recent work emphasizes explainability to foster trust, transparency, and user satisfaction (Burkart and Huber 2021). Early approaches integrated explainability into model architectures, such as explainable matrix factorization leveraging user similarities (Abdollahi and Nasraoui 2017), tree-enhanced embeddings (Wang et al. 2018), and attention-based neural models (Seo et al. 2017; Chen et al. 2018). More recent efforts have expanded explanation strategies, ranging from visual (Tsai and Brusilovsky 2019) and comparative methods (Le and Lauw 2021) to sentiment-oriented narratives (Chen, Yan, and Wang 2019), as well as traditional item-, content-, and popularity-based explanations (Sarwar et al. 2001; Pazzani and Billsus 2007; Yang 2016). These approaches highlight the importance of making recommendations more understandable, interactive, and socially aligned to support user trust and transparency across online communities.
Beyond these approaches, LLMs offer new opportunities for generating better explanations within the dynamic, user-driven landscapes (Mohanty 2024). By producing natural language explanations, LLMs can improve user satisfaction and trust, especially in domains where social context and user interaction are central (Chang, Harper, and Terveen 2016). For instance, LLMs have been used to create more detailed item descriptions (Acharya, Singh, and Onoe 2023) or to enable ChatGPT-based recommenders to generate user-friendly explanations (Deldjoo 2024). Recent research suggests that more contextualized, user-centered explanations could enhance trust and engagement (Silva et al. 2024; Lyu et al. 2024), reinforcing the value of LLM-driven strategies that align closely with the evolving social and informational needs of online communities.
2.2 User Perceptions of Recommendation Explanations
Users increasingly rely on recommender systems to navigate overwhelming volumes of information. Research grounded in the U&G framework highlights that individuals engage with media to fulfill specific needs, which can be categorized as affective, cognitive, and utilitarian motivations (Katz, Blumler, and Gurevitch 1973; Korhan and Ersoy 2016). These motivations directly influence consumption intentions, as users are more likely to engage with media or systems that effectively address their various needs. High-quality explanations within recommender systems can significantly influence user engagement, satisfaction, and decision-making, particularly in socially dynamic environments where these systems satisfy informational and social needs (Mandl et al. 2011; Jameson et al. 2015). For instance, explanations can enhance perceived usefulness and trust, aligning with motivation needs such as ease of use and task efficiency (Korhan and Ersoy 2016). Extant research also offers valuable insights into factors shaping user perceptions. For example, evidence shows that users may prefer human-generated recommendations, particularly for hedonic products (Wien and Peluso 2021). Similarly, explanations reflecting user feedback and personalization resonate with individuals who seek tailored, value-reinforcing content (Lu et al. 2023). Differences in user profiles further emphasize the importance of detailed and specific explanations, particularly for users with high cognitive engagement who value transparency and trust (Chatti et al. 2022).
Natural language explanations further enable systems to fulfill diverse needs, such as reducing cognitive effort, fostering trust, and enhancing community integration (Chang, Harper, and Terveen 2016). In parallel, researchers have developed evaluation metrics to address these user-centric demands beyond traditional measures like BLEU and ROUGE (Papineni et al. 2002; Lin 2004). User-centered metrics now consider dimensions such as effectiveness, persuasiveness, satisfaction, and trust, emphasizing user needs in complex, community-driven contexts (Zhou et al. 2021; Ehsan et al. 2019; Hoffman et al. 2023).
2.3 Research Gap
While the importance of recommendation explanations is widely acknowledged, research seldom examines how users interpret and respond to them. Furthermore, extant studies often rely on generic templates and have not fully explored the potential of LLM-generated contextualized explanations to address users’ cognitive, affective, and utilitarian needs (see Table 1).
To address this gap, our study provides the first comprehensive user evaluation of LLM-generated contextualized explanations through the lens of the U&G framework. By examining how these explanations resonate with users’ multifaceted needs and influence their intentions, we aim to offer deeper insights into designing recommendation explanations that are more aligned with user motivations.
Kunkel et al. Silva et al. Lubos et al. Our study Contextualized explanations Generic explanations Cognitive needs Affective needs Utilitarian needs Consumption intentions
Table 1: Comparison of research focus between previous studies and our work.
3 Method
3.1 Overview
We conducted an online between-subject experiment, where subjects ($N=759$) were randomly assigned to one of three experimental conditions: (1) recommendations with no explanations (NE), (2) recommendations with LLM-generated generic explanations (GE), and (3) recommendations with LLM-generated contextualized explanations (CE).
In the experiment, users first selected movies they had previously watched from a given set to establish the viewing history, based on which LLMs generated movie recommendations and explanations depending on the experimental condition. [^2] We used GPT-4 Turbo (version: gpt-4-1106-preview), which has knowledge of world events up to April 2023. For the experiment in the supplement, we used Claude 3.5 Sonnet. After receiving the corresponding explanation, users provided a series of ratings reflecting how well the explanation met underlying motivational needs, as well as their resulting behavioral intentions:
  • Cognitive needs: reflecting users’ desire for understanding and credibility in the recommendation process, assessed through Scrutability, Transparency, and Trust.
  • Affective needs: addressing the emotional and motivational dimensions of the user experience, evaluated through Persuasiveness and Satisfaction.
  • Utilitarian needs: focusing on practical, goal-oriented aspects of the recommendation, captured by Effectiveness, Efficiency, and Helpfulness.
  • Consumption intentions: reflect users’ behavior outcome, captured by UserIntent.
Dimensions composing the needs and intentions are established for assessing explanations in prior user studies (Balog and Radlinski 2020). We organize the dimensions under the U&G framework to depict users’ active pursuit of various needs from explanations. Given the statement for each dimension, users rated their agreement on a 7-point Likert scale from 1 (“Extremely Disagree”) to 7 (“Extremely Agree”). The higher-level intentions and needs are averaged on the specific dimensions (e.g., Cognitive is the average of Scrutability, Transparency, and Trust). A summary of the main variables and survey questions is in Table 2. After the survey, we randomly invited users to conduct follow-up interviews ($N=30$) to see their latent motivations across different explanations.
Variable Questions
Cognitive Needs
Transparency The text helps me to understand what the recommendation is based on.
Trust The text helps me to trust the recommendation.
Scrutability The text would allow me to give feedback on how well my preferences have been understood.
Affective Needs
Persuasiveness The text makes me want to watch this movie.
Satisfaction The text would improve how easy it is to pick a recommendation.
Utilirian Needs
Effectiveness The text helps me to determine how well I will like this movie.
Efficiency The text helps me to decide faster if I will like this movie.
Helpfulness How helpful do you find the recommendation?
Behvaiour Outcome
UserIntent How willing are you to watch this recommended movie?
Table 2: Summary of main variables and survey questions.
3.2 Subjects
Subjects were recruited from the U.S. via Cint. [^3] Cint: https://www.cint.com/ The online platform offers a built-in functionality for stratified sampling of subjects so that the sample is representative of the U.S. population. Subjects are remunerated with a fixed amount of around \$15 per hour.
As preregistered, we conducted a pilot study involving 50 subjects to assess the survey’s functionality and design. Informed by other user studies (Kang 2021; Lakens 2022), we performed a power analysis using G*Power 3.1 software (Faul et al. 2007, 2009). For multiple comparisons, we selected $F$-tests and chose a one-way ANOVA design with fixed effects (Lee and Lee 2018). We set the significance level ($\alpha$) to 0.05 and the power ($1-\beta$) to 0.95, with the adjusted effect size set at 0.1. The power analysis based on the pilot study results suggested that we need at least 750 subjects for the experiment.
Initially, 1401 subjects completed the survey. We discarded responses from subjects failing an initial attention check. Subsequently, following prior research (Pennycook, Cannon, and Rand 2018) and our pre-registration, a subset of subjects were excluded for three reasons: (i) subjects who self-reported to have answered randomly, (ii) failed the second attention check, (iii) responded excessively fast (i.e., the duration was less than around 6 seconds for responding to the user perceptions). These users still received remuneration after the exclusion. Eventually, due to our strict exclusion criteria, 759 of the 1,401 subjects qualified, exceeding our estimated sample size of 750 to ensure the results achieve the desired statistical power.
Among the 759 subjects, 387 are women, 370 are men, and two are non-binary. The sample has a mean age of 52.85 (SD = 16.01) and a range of 70 years (18 to 88). Most subjects have at least a high school diploma or equivalent.
3.3 Materials
We selected GPT-4 Turbo to generate movie recommendations for the main experiment. We conducted an additional experiment with Claude 3.5 Sonnet to validate the robustness of our study design and the generalizability of our findings (see supplementary materials). We precompiled all recommendations and explanations in a corpus to ensure that there was no variability in the outcomes across different LLM calls.
To create the corpus, we first selected 10 movies with diverse topics from the IMDb Top 250 list [^4] https://www.imdb.com/chart/top/ to compile a movie shortlist $\mathcal{M}$. These movies were chosen based on their high ratings or popularity. We ensured that the shortlist had broad coverage of user viewing preferences. Subsequently, we implemented a sequential recommendation strategy following the algorithm in Table 5 in Appendix A. With this strategy, sujects with the same viewing history and assigned conditions received the identical recommendation..
We followed previous practices (Giray 2023) when designing the prompts. Table 3 shows examples of contextualized and generic explanations. We can see a clear difference: while both explanations incorporate various elements (e.g., feature, popularity, similarity) to support the recommendations, the contextualized explanations naturally integrate the recommended movies with the user’s viewing history, explicitly highlighting how the user’s inputs shape the recommendation. Meanwhile, the generic explanations focus exclusively on the recommended movies. All generated movies were cross-checked with the IMDB dataset [^5] https://developer.imdb.com/non-commercial-datasets/ to ensure no hallucinations from the LLMs. Additionally, the corpus was validated by two experts with strong expertise in the field to ensure that all explanations used in the experiment were consistent and of fair quality.
System Prompt You are a movie recommendation assistant. Aim for clear, straightforward language that a broad audience can easily understand. Output each recommendation into three segments: The movie’s details, formatted as “Title - Release Year - Director”; A concise summary within 40 words that outlines the storyline; An explanation within 100 words. The explanation should focus on concrete aspects such as specific character arcs, notable plot developments, or memorable scenes.
User Prompt - CE For {Movie}, provide one recommendation for a user with a viewing history: {User viewing history}. Highlight how it directly appeals to the provided viewing history.
User Prompt - GE For {Movie}, write a recommendation explanation.
Table 3: Examples of contextualized and generic explanations generated by GPT-4 Turbo.
3.4 Procedure
During the experiment, we collected users’ viewing history and assigned them to one of our three conditions. Based on the condition assignment and viewing history, we generated the LLM-generated recommendations for them and collected their evaluations. The entire procedure was conducted online, consisting of the following steps:
Step 1. Users visit a welcome page where they provide informed consent to proceed with the experiment.
Step 2. Users read a short paragraph about the movie recommendation and respond to an attention check. Users who fail to pass are re-directed to the end of the survey. The attention check also serves to assess English proficiency.
Step 3. Users read about the purpose and functionality of a recommender system.
Step 4. Users select movies they have seen from the movie shortlist $\mathcal{M}$.
Step 5. Users provide the frequency of their movie consumption, their usage frequency, and the perceived helpfulness of recommender systems.
Step 6. Users are randomly assigned to one of our three conditions and shown corresponding recommendations.
Step 7. Following each recommendation, users express their perceptions across the dimensions of cognitive, affective, and utilitarian needs on a 7-point Likert scale. Users also state the recommendation’s helpfulness and express their intentions to consume the movie. Steps 6 and 7 are repeated for all movies selected by users from $\mathcal{M}$.
Step 8. Users go through a second set of attention checks.
Step 9. Users provide demographic information, i.e., age, gender, and education.
Step 10. Users participate in an honesty test where they are asked whether they answered all questions honestly, while being guaranteed that compensation would not be contingent on their answers.
Step 11. Users receive a debrief of the experiment and are informed that LLMs power all movie recommendations.
Afterward, we randomly reached out to $N=30$ users for the follow-up interviews. [^6] Detailed interview questions in Appendix B. We integrated their responses to help interpret the quantitative analysis results and investigate how explanations affect user motivational needs.
3.5 Analyses
We conducted analyses based on user perceptions in the survey. Each participant could provide up to 10 ratings, resulting in a total of 2,732 ratings from 759 participants in the main experiment (916 in NE, 898 in GE, and 918 in CE). We used hypothesis testing to identify differences in perceptions and consumption intentions across recommendation types, followed by mixed-linear regression to examine how user characteristics explain variation in both variables. Additionally, we conducted follow-up interviews to further interpret the quantitative results.
Hypothesis testing: We hypothesized that there are significant differences in the perceptions and consumption intentions between users receiving different recommendations. We conducted hypothesis testing on the four dependent variables of interest: Cognitive, Affective, Utilitarian, and UserIntent across the different conditions. Since the data did not meet the normality assumption (Lilliefors 1967), we used a Mann–Whitney test with Holm correction to identify pairwise differences between conditions (McHugh 2011). Later, we use $\mu$ to report the mean.
Mixed-effects linear regression. We used linear mixed-effects regression models to check further if any factors explained the variations in the three variables of interest. [^7] We also calculated the variance inflation factor (VIF) for all independent variables. All VIF values were below 5, indicating that multicollinearity should not bias the estimates (Akinwande, Dikko, and Samson 2015). We set NE as the reference condition so that our Treatment variables represent CE or GE. Given that each subject provides multiple ratings, we further included subject-specific random effects to control for between-subject heterogeneity (Linden and Hönekopp 2021). Additionally, to control for the habituation effect caused by repetitive measurements (Gaito 1961; Jankowski 2021), we included a control variable for the recommendation display order (i.e., DisplayOrder). Our final regression model is:
$\emph{Y}_{ij}$ $=\beta_{0}+\beta_{1}\emph{Treatment}_{j}+\beta_{2}\emph{Age}_{j}+\beta_{3}\emph{Gender}_{j}+\beta_{4}\emph{Education}_{j}$
$\quad+\beta_{5}\emph{MovieConsumption}_{j}+\beta_{6}\emph{RecSysFrequency}_{j}$ (1)
$\quad+\beta_{7}\emph{RecSysUtility}_{j}+\beta_{8}\emph{DisplayOrder}_{j}+u_{0j}+\epsilon_{ij}$
where ${Y}_{ij}$ denotes the dependent variables (i.e., Cognitive, Affective, Utilitarian, and UserIntent) for the $i$-th observation of the $j$-th subject, with intercepts $\beta_{0}$, subject-level random effects $u_{0j}$, coefficients $\beta_{1}$ through $\beta_{8}$, and an error term $\epsilon_{ij}$.
3.6 Ethical Considerations
We respect the privacy and agency of all subjects potentially impacted by this work and take steps to protect their privacy. The experiment design has obtained approval from the Ethics Commission at the University of Lausanne. All data were collected anonymously, and any personally identifiable features were removed. All steps during data collection and analyses followed standards for ethical research (Rivers and Lewis 2014). We also informed users that movie recommendations and explanations were generated by an LLM.
4 Results
Below, we state the results of the hypothesis tests. The regression results are also shown in Table 4.
4.1 Cognitive Needs
We observed a significant difference (Holm-adjusted $p<0.05$) in the distribution of the aggregated Cognitive measure (i.e., Scrutability, Transparency, Trust) between NE ($\mu=4.87$, $95\%$ CI = [4.78, 4.95]) and CE ($\mu=5.22$, $95\%$ CI = [5.14, 5.30]). We also found a significant difference between GE and CE (Holm-adjusted $p<0.05$), with CE surpassing GE ($\mu=4.90$, $95\%$ CI = [4.81, 4.99]). However, there was no statistical difference between NE and GE (Holm-adjusted $p>0.05$), suggesting that generic explanations do not differ markedly from no explanations in meeting users’ cognitive needs.
In the mixed-effects regression, both RecSysUtility ($\beta=0.319$, $p<0.05$) and MovieConsumption ($\beta=0.149$, $p<0.05$) demonstrated positive, significant effects on Cognitive, indicating that participants who perceive the recommender system as more useful and who watch movies more frequently tend to rate higher in Cognitive.
In the follow-up interviews, several participants expressed increased trust and clarity when receiving the contextualized explanations. One participant noted: “When it told me the rationale for suggesting the movie, I could decide whether or not to watch it.” Another participant stated: “Including the movie I had watched and comparing the similarities made me trust the recommendation more.” Interestingly, some users also noted that references to movies they disliked helped them filter out unwanted content, improving their sense of control and overall transparency.
These quantitative and qualitative findings suggest that contextualized explanations, by explicitly linking a user’s past and present preferences, can bolster trust and understanding in a way that generic or absent explanations do not. The more personal and concrete presentation of information serves users’ cognitive needs by explaining why some recommendations might be a good match, and others might not align with their tastes, ultimately leading to a deeper sense of scrutiny, transparency, and trust.
Figure 2: Cognitive ratings by explanation type (shown variable: Cognitive). Significance levels between conditions are indicated: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$.
4.2 Affective Needs
We found no overall difference in the aggregated Affective measure (i.e., Persuasiveness, Satisfaction) among NE ($\mu=4.97$, $95\%$ CI = [4.88, 5.05]), GE ($\mu=4.92$, $95\%$ CI = [4.83, 5.01]), and CE ($\mu=5.02$, $95\%$ CI = [4.93, 5.11]). However, examining each subdimension showed that Persuasiveness differed significantly between GE and CE (Holm-adjusted $p<0.05$), indicating that contextualized explanations can make recommendations more compelling than generic explanations, even if this advantage does not fully carry over to overall perceived affective needs.
The mixed-effects regression highlights two user-engagement variables, RecSysUtility ($\beta=0.311$, $p<0.05$) and MovieConsumption ($\beta=0.167$, $p<0.05$), as positive, significant predictors of affective responses. This suggests that users who view the recommender system as helpful and frequently watch movies are more likely to feel emotionally drawn in, regardless of the explanation style.
Our interviews offer deeper insights into how these affective needs take shape. Several participants emphasized that a good explanation heightened their eagerness to watch the movie: “They made me want to watch them more than I would have before with their explanations.” Others found “overly pushy” or “over the top” descriptions off-putting, highlighting the delicate balance between building excitement and overselling the film. Some appreciated specific references to actors or plot details, while others disliked spoiler-like content, favoring a brief, suggestive teaser. These varying reactions prove that users’ emotional engagement hinges on contextual cues and personal taste, mood, and openness to promotional language.
These findings imply that satisfying affective needs, such as feeling entertained, inspired, or emotionally “in tune”, goes beyond simply adding contextual detail. Users’ enjoyment levels are tightly bound to their existing trust in the system, general interest in watching movies, and thresholds for promotional or spoiler content. Consequently, while contextualized explanations enhance persuasiveness for some users, effectively catering to affective motivations requires a more nuanced approach that recognizes diverse emotional triggers and preferences.
Figure 3: Affective ratings by explanation type (shown variable: Affective).
4.3 Utilitarian Needs
There is no overall difference in the aggregated Utilitarian measure (i.e., Effectiveness, Efficiency, Helpfulness) across NE ($\mu=5.24$, $95\%$ CI = [5.16, 5.32]), GE ($\mu=5.27$, $95\%$ CI = [5.19, 5.36]), and CE ($\mu=5.27$, $95\%$ CI = [5.18, 5.35]). However, a closer examination at Helpfulness revealed that NE ($\mu=5.07$, $95\%$ CI = [4.98, 5.17]) versus CE ($\mu=5.23$, $95\%$ CI = [5.13, 5.33]) was significant (Holm-adjusted $p<0.05$), suggesting that contextualized explanations can offer an advantage over having no explanation in practice by helping users assess whether a movie meets their needs. By contrast, there was no difference between GE and CE in Helpfulness, indicating that once any form of explanation is provided, adding contextual details may not further enhance users’ perceived practical value in selecting movies.
The mixed-effects regression showed RecSysUtility ($\beta=0.292,p<0.05$) and MovieConsumption ($\beta=0.181,p<0.05$) both positively influenced Utilitarian, again implying that users who closely engage with recommender systems and movies, tend to see greater practical value in the recommendations regardless of explanation style. Female participants also reported slightly higher utilitarian scores (versus male, $p<0.05$).
Our interview provides further evidence. Several participants appreciated having clear, concise details about the plot or genre, emphasizing how these insights save time in deciding whether a movie fits their tastes: “The more it detailed the movie, the more it helped me decide.” Some also suggested that referencing both liked and disliked movies aided them in filtering out poor matches and reinforcing good ones: “I was able to assess whether I wanted to watch it, based on what I liked before, and also on what I knew I didn’t enjoy.” Others felt the explanations were too vague or enthusiastic, hindering their ability to gauge true quality.
Although contextualized explanations may slightly outperform no explanations in offering helpful guidance, overall utilitarian gains do not appear large when considering Effectiveness or Efficiency. These findings suggest that users’ utilitarian needs stem less from how the explanation is framed and more from whether the system is already perceived as useful and aligns with existing consumption habits. Nonetheless, highlighting specific points of similarity or dissimilarity can provide an additional benefit in helping users quickly identify which recommendations are worth pursuing and which to skip.
Figure 4: Utilitarian ratings by explanation type (shown variable: Utilitarian).
4.4 Consumption Intentions
Knowing users’ consumption intentions (measured by UserIntent) helps us understand how explanations translate into actual viewing decisions. Pairwise comparisons confirm that CE ($\mu=4.70$, $95\%$ CI = [4.59, 4.81]) significantly differs from GE ($\mu=4.31$, $95\%$ CI = [4.18, 4.43]) (Holm-adjusted $p<0.05$). The contrasts between NE ($\mu=4.53$, $95\%$ CI = [4.42, 4.65]) and CE, as well as NE and GE, are also significant (Holm-adjusted $p<0.05$). This pattern suggests that the existence of explanations and contextualization fosters greater motivation to watch a recommended movie.
Mixed-effects regression indicates that CE increases consumption intentions by 0.240 points on average ($\beta=0.240,p<0.05$) compared to GE. RecSysUtility ($\beta=0.336,p<0.05$) and MovieConsumption ($\beta=0.143,p<0.05$) both increase the likelihood of watching the recommended title. Meanwhile, older participants ($\beta=-0.139,p<0.05$) and female users ($\beta=-0.227,p<0.05$) expressed slightly lower consumption intentions.
The qualitative interview responses underscore how alignment with personal taste can motivate users to watch the recommended movies. Several participants mentioned that recognizable themes, plot twists, or well-known actors greatly influenced their choices. One participant explained: “An interesting twist or variation from an expected plot development intrigues me, as does the presence of an actor I like.” Meanwhile, others highlighted that the mention of disliked or irrelevant elements quickly led them to dismiss a recommendation: “Some movies I simply will not watch … if it’s not my kind of genre, no explanation can convince me.” In some instances, participants weighed social factors such as the opinions of friends and family members, underscoring how shared viewing decisions can override individual preferences if, for example, a partner or group prefers a different type of film.
These results imply that while contextualized explanations provide an edge over generic ones, individual beliefs about system helpfulness and prior consumption habits remain critical in determining whether users actually watch a recommended movie. Explaining “why this movie” helps, but so does the user’s fundamental perception of the recommender’s reliability and personal engagement with the medium. Ultimately, consumption intentions emerge from a synthesis of explanation quality, personal taste, and social or experiential factors, illustrating that users weigh both the persuasive power of contextualization and their unique circumstances.
Figure 5: Consumption intentions by explanation type (shown variable: UserIntent). Significance levels between conditions are indicated: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$.
Cognitive Affective Utilitarian UserIntent
Intercept
4.870***
4.906***
5.084***
4.642***
Treatment (GE)
0.030
--0.010
0.082
--0.180
Treatment (CE)
0.366***
0.125
0.118
0.240*
Gender (Female)
0.094
0.134
0.284***
--0.227*
Age
--0.061
--0.076
0.025
--0.139*
Education
0.049
0.047
0.019
0.060
RecSysUtility
0.319***
0.311***
0.292***
0.336***
RecSysFrequency
0.013
--0.004
--0.043
0.106
MovieConsumption
0.149***
0.167***
0.181***
0.143*
DisplaypOrder
--0.018
--0.046
--0.031
--0.022
Significance levels: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$
Table 4: Mixed-effect linear regression results. Coefficients are shown with standard errors in parentheses.
5 Supplementary Experiments
To supplement the main experiment, we expanded the experiment to a larger scope by leveraging another LLM (Claude 3.5 Sonnet).
Given that GPT-4 Turbo is one of the most powerful LLMs, we tested whether the superior performance of contextualization could be generalized to other LLMs. We generated another recommendation corpus using Claude 3.5 Sonnet and replicated the experiments on the U.S. samples ($N=140$, 71 are women, 68 are men, and one is non-binary). The results are aligned with our main findings, validating the advantages of LLMs in generating contextualized explanations in terms of meeting users’ motivational needs and increasing their consumption intent. The results and interpretations are provided in Appendix C.
In summary, the supplementary experiment yielded results that supported the main experiment’s findings and conclusions, thereby validating the robustness of our study design and the generalizability of the findings.
6 Discussion
6.1 Implications of Findings
Recommender systems play a central role in shaping user decisions, particularly influential in the context of web and social media. Given the far-reaching implications of these systems, there is growing demand from users and regulators for transparency, emotional connection, or efficiency in decision-making. This highlights the value of high-quality and context-aware recommendation explanations. However, many existing recommender systems do not or fail to provide explanations that address users’ diverse needs (Li, Zhang, and Chen 2020; Lu et al. 2023).
Our study demonstrates that LLM-generated contextualized explanations can overcome these limitations by satisfying users’ cognitive needs, ultimately increasing their consumption intentions. Also, our findings reveal the insufficiency of LLM-generated explanations in meeting users’ utilitarian and affective needs, which raises concerns about the proper design and implications of LLM-generated personalized content.
Benefits of contextualization. Our results show that contextualized explanations significantly satisfy users’ needs and their intention to watch the recommended movies. This is particularly evident in users’ cognitive needs (Scrutability, Transparency, Trust), where contextualization not only promotes a clearer understanding of why specific titles are recommended but also encourages users to scrutinize recommendations more carefully. The increase in trust and perceived transparency shows the importance of tailoring explanations to individual needs. Offering personally relevant and meaningful recommendations can encourage users to act on these suggestions, especially in entertainment and e-commerce settings, where a personalized touch can be key to customer retention.
LLMs in the recommendation. As LLMs evolve, their capacity to generate richer, more personalized explanations will continue to grow (Feuerriegel et al. 2025). This capability is particularly relevant for companies like Amazon and Netflix, where personalization is a competitive differentiator (Burroughs 2019). Our findings highlight how integrating LLMs into recommender systems can satisfy users’ cognitive needs and increase their consumption intention. Beyond driving engagement, context-aware explanations may help companies fulfill emerging regulatory requirements, such as the EU’s AI Act, which calls for transparency and the “right to explanation” in AI-driven decisions (Veale and Borgesius 2021; You and Suh 2024).
System utility and positive feedback loops. Our study suggests that a user’s active and positive engagement with the recommender system amplifies the benefits of contextualization. When participants view the system as helpful and relevant to their interests, context-aware explanations further boost their ratings of recommendation quality, trust, and helpfulness. This synergy can create a positive feedback loop: as users find the system increasingly beneficial, they engage more deeply, and contextualized explanations, in turn, reinforce these favorable perceptions. Moreover, frequent movie watchers also exhibit higher cognitive, affective, and utilitarian evaluations. Increasing user engagement through personalized interfaces, easy-to-use features, or curated content could, therefore, magnify the impact of contextualized explanations (Lacker and Way 2024). Companies seeking to cultivate sustained user relationships can harness this dynamic by combining context-rich recommendation details with user education or incentives that further encourage exploration and use of the system.
Avoiding redundant information. An interesting insight from the experiment is that some participants found overly “pushy” or “over the top” descriptions off-putting. In contrast, others appreciated detailed explanations. Although contextualized explanations can meet various user needs, avoiding irrelevant or excessive information that increases cognitive load is critical. Our findings suggest that, for specific users, no explanation is preferable to offering a generic one. This reflects a balancing act: well-crafted, context-specific explanations can significantly improve user perceptions, but lengthy or poorly targeted explanations can become a distraction.
In industries such as travel and finance, where decisions can be complex, selective use of contextualized explanations could help users make informed choices without overwhelming them. By highlighting only the most pertinent details, companies can streamline the decision-making process, thus promoting higher engagement and satisfaction.
6.2 Limitations
While our study offers valuable insights into user perceptions of LLM-generated explanations for movie recommendations, several limitations should be noted.
First, we primarily recruited English-speaking users. Future research should include more diverse populations to strengthen the generalizability of our findings.
Second, the effectiveness of the explanations may have been influenced by the performance of the underlying LLMs. Although we used state-of-the-art models (GPT-4 Turbo and Claude 3.5 Sonnet), we should notice that future advancements in LLMs could further shape or change user perceptions.
Third, our distinction between contextualized and generic explanations may oversimplify the explanation types. Future studies should explore a broader variety of explanation strategies to better understand their impact in different contexts and on diverse user groups.
Lastly, our experimental design did not fully capture the dynamic real-world user interactions with recommender systems. User perceptions might change after consuming the recommended movies, which is not covered by this study. This highlights the necessity of conducting the user study in real scenarios.
Despite these limitations, our study provides a strong foundation for further research into the role of LLM-generated explanations in recommender systems, emphasizing the importance of contextualization in improving user experience and engagement.
7 Conclusion
Our findings show how and why recommendation explanations meet cognitive, affective, and utilitarian motivations under the U&G perspective. Our study confirms that LLM-generated contextualized explanations effectively address users’ cognitive needs and increase their consumption intentions. Notably, recommender systems should also avoid overwhelming users with irrelevant information that can hinder decision-making or erode trust. Future recommender system designs that integrate contextualized LLM-generated explanations can more effectively serve a wide range of user motivations, ultimately improving user satisfaction and leading to actionable engagement with the recommended content.
Appendix A Recommendation Corpus Generation
Recommendation Corpus Generation
Input: Movie shortlist $\mathcal{M}$
Output: Corpus $\mathcal{C}$ containing recommendations and explanations
1 for each $m_{i}\in\mathcal{M}$ do
   $m_{i}$: Represents a single movie in the shortlist $\mathcal{M}$
2    for $k=1$ to 4 do
   $k$: The recommendation order (1st, 2nd, 3rd, or 4th)
3     if $k=1$ then
4      Generate $r_{i1}$ from $m_{i}$
     $r_{ik}$: The $k$-th recommendation for the $i$-th movie
5     else
6      Generate $r_{ik}$ from $\{m_{i},r_{i1},\dots,r_{i(k-1)}\}$
     Previous recommendations are incorporated to generate $r_{ik}$
7     end if
8     Generate $e_{ikc}$ and $e_{ikg}$ for $r_{ik}$
    $e_{ikc}$: Contextualized Explanation for $r_{ik}$
    $e_{ikg}$: Generic Explanation for $r_{ik}$
9     Save $r_{ik}$, $e_{ikc}$, and $e_{ikg}$ in $\mathcal{C}$
    Corpus $\mathcal{C}$ stores the recommendations and their explanations
10    end for
11 end for
Table 5: Recommendation corpus preparation steps. The corpus $\mathcal{C}$ contains recommendations $r_{ik}$ for each movie $m_{i}\in\mathcal{M}$, with contextualized explanations ($e_{ikc}$) and generic explanations ($e_{ikg}$) for each recommendation.
Appendix B Interview Questions
  • “What made you trust or doubt the recommendation after seeing the explanation?”
  • “How did the explanation affect your overall feeling about the movie? Why did it make you feel that way?”
  • “How did the explanation assist (or fail to assist) you in understanding if the movie fit your interests or needs?”
  • “After reading the explanation, what influenced you most in deciding whether to watch the movie or not?”
  • “How do you think your profile, such as your demographics or viewing habits, influenced how you felt about the explanation you received?”
Appendix C Supplementary Experiment (with Claude 3.5 Sonnet)
We replicated the experiment on a smaller U.S. sample ($N=140$) using Claude 3.5 Sonnet. The results align with our main study, showing that contextualization improves overall explanation quality, particularly in terms of persuasiveness, transparency, and trust. Moreover, contextualized explanations increase users’ consumption intentions. Additionally, the positive effect of RecSysUtility is maintained, even with another underlying LLM instead of GPT-4 Turbo. While some minor discrepancies emerged between this supplementary experiment and the main experiment, the core findings remain consistent. This supports the generalizability of our study’s conclusions. Figure 6, Figure 7, Figure 8, and Figure 9 show the user ratings of motivational needs and consumption intentions in this experiment.
Figure 6: Cognitive ratings by explanation type (shown variable: Cognitive). Significance levels between conditions are indicated: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$.
Figure 7: Affective ratings by explanation type (shown variable: Affective).
Figure 8: Utilitarian ratings by explanation type (shown variable: Utilitarian).
Figure 9: Consumption intentions by explanation type (shown variable: UserIntent). Significance levels between conditions are indicated: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$.
Cognitive Affective Utilitarian UserIntent
Intercept
4.823***
4.807***
5.103***
4.559***
Treatment (GE)
0.097
0.112
--0.055
0.101
Treatment (CE)
0.509*
0.534*
0.331
0.730*
Gender (Female)
--0.126
--0.228
0.064
--0.499
Age
0.269*
0.234*
0.274*
0.145
Education
--0.006
--0.040
--0.052
0.053
RecSysUtility
0.217*
0.155
0.268
0.418*
RecSysFrequency
0.156
0.146
0.076
0.030
MovieConsumption
0.044
--0.018
--0.045
--0.148
Significance levels: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$
Table 6: Mixed-effects linear regression results for the four dependent variables: Cognitive, Affective, Utilitarian, and UserIntent. Coefficients are shown with standard errors in parentheses.